72

A 2.0g sample of steam at 100C loses 1140 calories of heat. What is theresulting temperature of the sample?
Answer #1

Given Q = 1140cal = 1140*4.18J = 4765Now Q = (2*10-3) L +msΔT4765=(2*10-3)[(2.26*106) + (4186) ( 100 - Tf)2382 *103=(2.26*106)+ (4186) ( 100)-4186 *TThe temp = 70.85
answered by: Kayka